OpenStudy (anonymous):

Line CD passes through points (0, 2) and (4, 6). Which equation represents line CD? y = 2x – 2 y = 2x + 2 y = x + 2 y = x – 2

OpenStudy (anonymous):

To find the slope: (y1-y2)/x1-x2) (6-2)/(4-0) slope = 1 use one of the points' coordinates to find b 2=1(0)+b b=2 C